### v1

# DPCL-Diff: The Temporal Knowledge Graph Reasoning based on Graph Node Diffusion Model with Dual-Domain Periodic Contrastive Learning

Abstract
Temporal knowledge graph (TKG) reasoning that infers future missing facts is an essential and challenging task. Predicting future events typically relies on closely related historical facts, yielding more accurate results for repetitive or periodic events. However, for future events with sparse historical interactions, the effectiveness of this method, which focuses on leveraging high-frequency historical information, diminishes. Recently, the capabilities of diffusion models in image generation have opened new opportunities for TKG reasoning. Therefore, we propose a graph node diffusion model with dual-domain periodic contrastive learning (DPCL-Diff). Graph node diffusion model (GNDiff) introduces noise into sparsely related events to simulate new events, generating high-quality data that better conforms to the actual distribution. This generative mechanism significantly enhances the model’s ability to reason about new events. Additionally, the dual-domain periodic contrastive learning (DPCL) maps periodic and non-periodic event entities to Poincaré and Euclidean spaces, leveraging their characteristics to distinguish similar periodic events effectively. Experimental results on four public datasets demonstrate that DPCL-Diff significantly outperforms state-of-the-art TKG models in event prediction, demonstrating our approach’s effectiveness. This study also investigates the combined effectiveness of GNDiff and DPCL in TKG tasks.
DPCL-Diff: The Temporal Knowledge Graph Reasoning based on Graph Node Diffusion Model with Dual-Domain Periodic Contrastive Learning
Yukun Cao1, Lisheng Wang2, Luobing Huang3thanks: 123The authors are affiliated with the School of Computer Science and Technology, Shanghai University of Electric Power (SUEP), Shanghai 201306, China thanks: 1First author (email: marilyn_cao@163.com)thanks: 2Corresponding author (email: 2207992575@mail.shiep.edu.cn)
1 Introduction
Knowledge graphs (KGs) store a large amount of information about the natural world and have shown great success in many downstream applications, such as natural language processing Hu et al. (2022); Wang et al. (2020), recommendation system Xuan et al. (2023); Yu et al. (2022), and information retrieval Zhou et al. (2022). Traditional KGs typically only contain static snapshots of facts, integrating facts (also known as events) in the form of static relational triples $(s,r,o)$, where $s$ and $o$ represent the subject and object entities, respectively, and $r$ represents the relationship type. However, in the real world, knowledge evolves and constantly exhibits complex temporal dynamics Wang et al. (2017); Yin and Cui (2016), which has inspired the construction and application of Temporal Knowledge Graph (TKG). TKG extends the previous static relational triples $(s,r,o)$ to quaternions $(s,r,o,t)$ with timestamps. Thus, the TKG consists of multiple snapshots where facts in the same snapshot appear simultaneously.
Existing studies have identified two main types of new events in TKG: periodic events that will recur Liang et al. (2023); Li et al. (2021), and events that have not occurred before but may occur in the future Xu et al. (2023). For these new events, TKG reasoning offers fresh perspectives and insights for many downstream applications, such as event prediction Zhou et al. (2023), political decision making Zhang and Zhou (2022), dialogue generation Hou et al. (2024), and text generation Li et al. (2022b). These applications have greatly stimulated intense interest in TKG. In this work, we focus on predicting new facts in future time, a task known as graph extrapolation Chen et al. (2023); Li et al. (2021). Our goal is to predict the missing entities in the query $(s,p,?,t)$ for future timestamps $t$ that still need to be observed in the training set.
Most current work models the structural and temporal characteristics of TKG to capture the specific relationships and temporal dependencies between different events for future event prediction.Many studies Liang et al. (2023); Xu et al. (2023); Li et al. (2021)can predict repetitive or periodic events by referring to the known historical events and distinguishing the different effects of periodic and non-periodic events on reasoning tasks through contrastive learning. However, in actual reasoning tasks, some periodic events share the same head entities and relations, differing only in their tail entities. This results in overly similar representations, making it difficult to distinguish them during reasoning. Additionally, In event-based TKG, new events that have never occurred account for about 40% Jäger (2018). Due to the sparse traces of temporal interactions throughout the timeline of these new events, it is impossible to use reasoning methods based on high-frequency periodic events, resulting in poor reasoning performance for these types of events.
To address the above problem, we propose a diffusion model-based dual-domain periodic contrast learning for timing knowledge graph reasoning (DPCL-Diff), which employs a Graph Node Diffusion model (GNDiff) to reason about new events. GNDiff tackles the issue of sparse interaction traces for new events on the timeline by introducing noise into existing correlated sparse events of the new entities. This diffusion mechanism simulates the real-world occurrence of new events and is gradually refined during the denoising process, guided by pre-trained language models (PLMs) that have been trained on extensive corpora and possess a foundational understanding of background knowledge. Consequently, this approach generates high-quality data that closely aligns with the actual distribution. The diffusion model Ho et al. (2020); Song et al. (2020) is a cutting-edge generative framework with demonstrated powerful capabilities in both image Ramesh et al. (2022); Rombach et al. (2022); Saharia et al. (2022)and text generation Ou and Jian (2024); Gong et al. (2022); Strudel et al. (2022). This provides more correlation information for inferring new events and dramatically improves the ability to reason about new events in TKG. For event inference tasks containing a large amount of temporal interaction information, we propose a novel dual-domain periodical contrast learning (DPCL) method. The periodic and non-periodic event entities related to the current event are mapped into the Poincaré and Euclidean spaces for contrastive learning, respectively. The Poincaré space property is utilized to distinguish the similar periodic event entities better.
The main contributions of this thesis are summarized below:
• We propose a TKG model called DPCL-Diff. DPCL-Diff can reason not only about periodic events but also about new events through a diffusion generation mechanism.
• GNDiff introduces a novel approach in TKG. GNDiff leverages pre-trained language models (PLMs) to conduct non-Markov decision chain-based diffusion for new events, generating high-quality data and effectively addressing the challenge of sparse interaction traces.
• We propose the DPCL method, where periodic and non-periodic event entities are mapped into the Poincaré space and Euclidean space to better distinguish similar periodic event entities.
• We conducted experiments on four public datasets. The results show that DPCL-Diff outperforms the state-of-the-art TKG model in event prediction. Additionally, we purposefully explore the effectiveness of GNDiff and DPCL in the TKG task.
Figure 1: Overall architecture of DPCL-Diff. The first part is GNDiff, and the second part is DPCL.
2 Related Work
2.1 Diffusion Model
Diffusion models Ho et al. (2020) are latent variable models designed for continuous data, achieving state-of-the-art sample quality in image and audio generation. To handle discrete data, prior work explored text diffusion models in discrete state spaces and defined diffusion processes for such data Li et al. (2024); Lezama et al. (2022); Austin et al. (2021). Li et al. (2022a) focuses on continuous diffusion modelling for text, enabling gradient-based methods to control generation by introducing continuous latent representations. Ou and Jian (2024) combines discrete diffusion models with denoising techniques, significantly improving text generation quality and diversity in unconditional and conditional tasks.
2.2 Temporal Knowledge Graph Reasoning
TKG reasoning can be approached through interpolation or extrapolation Chen et al. (2023); Li et al. (2021). Interpolation addresses missing events between $t_{0}$ to $t_{n}$,while extrapolation predicts events beyond $t_{n}$. Interpolation models like TeAST Li et al. (2023), Re-Temp Wang et al. (2023), and TempCaps Fu et al. (2022) infer missing relationships but cannot predict future events. Extrapolation methods focus on future event prediction. Know-Evolve Trivedi et al. (2017), the first to learn nonlinear evolutionary entity embeddings, struggles with long-term dependencies. CyGNet Zhu et al. (2021) uses replication generation to identify frequent repetitive events, while CENET Xu et al. (2023) builds on this with contrastive learning to identify highly correlated entities. TARGAN Xie et al. (2023) enhances long-term modeling using attentional mechanisms. Other models, including DynaQ Liu et al. (2024a), EvoGNN Pareja et al. (2020), and TempoR Wu et al. (2020) have improved prediction but still face challenges in jointly capturing temporal evolution, semantic dependencies, and scalability.
3 Method
As shown in Fig. 1, for novel events, the DPCL-Diff model leverages GNDiff to inject noise into the sparse correlated events of the new event entity. This approach simulates the real-world mechanism of new events, generating numerous high-quality data samples. For other events, the DPCL method identifies highly correlated entities through comparative learning and better distinguishes similar periodic event entities using Poincaré space properties. In the following few sections, we describe our proposed approach in detail.
3.1 Preliminaries
TKG $G=\{G_{0},G_{1},...,G_{t},...,G_{T}\}$ is a sequence of KGs, each $G_{t}=\{\mathcal{E},\mathcal{R},\mathcal{T}\}$ contains facts that occur at timestamp $t$. Let $\mathcal{E}$, $\mathcal{R}$ and $\mathcal{T}$ denote a finite set of entities, relation types, and timestamps, respectively. Each fact is a quadruple $(s,r,o,t)$ , where $s\in\mathcal{E}$ is the subject entity, $o\in\mathcal{E}$ is the object entity, and $r\in\mathcal{R}$ is the relation at timestamp $t$ between $s$ and $o$ that occurs, the dimension is $d$. ||×$\mathcal{E}\in\mathbb{R}^{|\mathcal{E}|\times d}$ is the embedding of all entities, and its rows represent the embedding vectors of an entity, such as $s$ and $o$. Similarly,||×$\mathcal{R}\in\mathbb{R}^{|\mathcal{E}|\times d}$ is the embedding of all relation types. The TKG inference task aims to predict missing objects by answering a query $q=(s,r,?,t)$ using a given history of TKG.
For a given query $q=(s,r,?,t)$, we define the set of periodic events $\mathcal{Q}_{t}^{s,r}$ as the set of events that have the same head entity relationship with the query event, and the corresponding set of periodic event entities as $\mathcal{H}_{t}^{s,r}$ in the following equations:
$\mathcal{Q}_{t}^{s,r}=\bigcup_{k<t}\{(s,r,o,k)\in\mathcal{G}_{k}\},$ (1)
$\mathcal{H}_{t}^{s,r}=\{o|(s,r,o,k)\in\mathcal{Q}_{t}^{s,r}\}.$ (2)
Entities outside $\mathcal{H}_{t}^{s,r}$ are non-periodic. If an event $(s,r,o,t)$ does not exist in $\mathcal{Q}_{t}^{s,r}$, it is a new event that has never occurred and is handled by the GNDiff. For periodic event reasoning, DPCL is primarily used to predict target entities. The following sections detail these modules.
3.2 Graph Node Diffusion Model
In this section, we introduce GNDiff, showcasing the potential of diffusion models for TKG applications. While previous work has effectively explored the diffusion of discrete textual data, applying diffusion models to the discrete graph nodes in TKGs can be similarly effective. Traditional reasoning methods, which rely on historical data to learn implicit associations, need help to reason about new events due to the sparse historical data associated with them. In contrast, GNDiff employs a diffusion model to simulate the event generation mechanism, producing many high-quality samples for new event inference.
GNDiff consists of a forward diffusion process and a reverse diffusion process. For each quadruple $(s,r,o,t)$, after obtaining the vector representations of $s$, $r$, and $o$, these vectors are concatenated into a single vector, which serves as the input to the graph node diffusion model. We then obtain the sample data $x_{0}\sim q(x_{0})$. Since TKGs involve discrete data, each element of $x_{t}$ is a discrete random variable with $K$ categories. For graph node, $K=\left|\mathcal{V}\right|$ represents the number of node types. Representing $x_{t}$ as a stack of heat vectors, the process of adding noise is denoted as shown in Equation 3:
$q(x_{t}\mid x_{t-1})=\mathrm{Cat}(x_{t};p=x_{t-1}Q_{t}),$ (3)
where $\mathrm{Cat}(\cdot)$ denotes that the noise addition process is based on the category distribution, and ${Q}_{t}$ is used to control the transition matrix of the node from one state to another:
$[Q_{t}]_{i,j}=q(x_{t}=j\mid x_{t-1}=i),$ (4)
The conversion relationship is shown in Equation 5:
${q}(x_{t-1}\mid x_{t},x_{0})=\frac{q(x_{t}\mid x_{t-1},x_{0})q(x_{t-1}\mid x_{0})}{q(x_{t}\mid x_{0})}$ (5)
$=\mathrm{Cat}\Bigg{(}x_{t-1};p=\frac{x_{t}Q_{t}^{\top}\odot x_{0}\overline{Q}_{t-1}}{x_{0}\overline{Q}_{t}x_{t}^{\top}}\Bigg{)},$
where $\overline{Q}_{t}=Q_{1}Q_{2}\cdots Q_{t}$. Note that $\odot$ denotes elementwise multiplication, and division is performed rowwise. To be combined with pre-trained denoising language models, we incorporate an absorbing state, allowing for the generation of samples during denoising. We use the masking mechanism to simulate the possible inactive state of the node. The transition matrix adjustment process is shown in Equation 6:
$[Q_{t}]_{i,j}=\begin{cases}1&\mathrm{if}\quad i=j=[\mathrm{M}],\\ \beta_{t}&\mathrm{if}\quad j=[\mathrm{M}],\ \mathrm{i}\neq[\mathrm{M}],\\ 1-\beta_{t}&\mathrm{if}\quad i=j\neq[\mathrm{M}],\end{cases}$ (6)
Here, $[\mathrm{M}]$ denotes a node in a masked state (an inactive node), allowing nodes to transition naturally in and out of this state as the graph evolves, eventually converging to a smooth distribution $q(x_{T})$. This approach concentrates all probability mass on a sequence labelled with a mask. By using $\overline{Q}_{t}=Q_{1}Q_{2}\cdots Q_{t}$ , the transfer relation for multiple time steps can be derived. Samples are then generated through a backward diffusion process:
$p_{\theta}({x}_{0:T})=p({x}_{T})\prod_{t=1}^{T}p_{\theta}({x}_{t-1}\mid{x}_{t},t),$ (7)
We optimize the inverse process by minimizing the variational lower bound to maximize the log-likelihood of the original data. This approach derives the optimization objective $p_{\theta}({x}_{0:T})$ and controls the optimization through $\mathcal{L}_{\mathrm{diff}}$:
$\mathcal{L}_{\mathrm{diff}}=\mathbb{E}_{q}[\mathrm{D}_{\mathrm{KL}}(q(x_{T}\mid x_{0})\parallel p_{\theta}(x_{T}))]+$ (8)
$\mathbb{E}_{q}[\sum_{t=2}^{T}\mathrm{D}_{\mathrm{KL}}(q({x}_{t-1}|{x}_{t},{x}_{0})\parallel p_{\theta}({x}_{t-1}|{x}_{t},t))]$
$-\log p_{\theta}(x_{0}\mid x_{1}),$
where ${E}_{q}$ denotes the expectation under a noisy data distribution. $\mathrm{D}_{\mathrm{KL}}$ stands for Kullback-Leibler divergence.
3.3 Dual-Domain Periodic Contrastive Learning
Inspired by Xu et al. (2023), we propose a dual-domain periodic contrastive learning method. Unlike the original approach, we use a dual-domain mapping strategy because we observed that specific queries involve multiple similar periodic events that are challenging to distinguish. To address this issue, we map entities from periodic and non-periodic events to Poincaré and Euclidean space, respectively, to identify highly correlated entities within each domain. By leveraging the spatial characteristics of Poincaré space, we effectively differentiate similar periodic entities.
During data preprocessing, we analyze entity frequencies for a given query $q=(s,r,?,t)$, as shown in Equation 13:
$\mathbf{Z}_{t}^{s,r}(o)=\lambda\cdot(\Phi_{\beta>0}-\Phi_{\beta=0}),$ (9)
$\beta=\sum_{(s,r,o_{i},k)\in G_{k}}(o_{i}=o),$ (10)
where the value of each entity frequency is constrained by the hyperparameter $\lambda$. $\Phi_{\beta}$ is an indicator function that returns 1 if $\beta$ is true and 0 otherwise. $\mathbf{Z}_{t}^{s,r}(o)>0$ indicates that quadruple $(s,r,o,t_{k})$ is a periodic event bound to $s$ and $r$; otherwise,it is a non-periodic event. DPCL then learns dependencies in periodic and non-periodic events based on the input $\mathbf{Z}_{t}^{s,r}(o)$. The distance between entities in Poincaré space is used as the final similarity score between an entity $o_{i}\in\mathcal{E}$ and a query $q$. The formula for calculating the Poincaré distance is provided as follows:
$\mathbf{S}_{p}^{s,r}(o_{i})=tanh(\mathbf{W}_{p}{s}\oplus r+\mathbf{b}_{p})\mathbf{E}^{T}+\mathbf{Z}_{t}^{s,r}+\mathbf{d}_{p}(s,o_{i}),$ (11)
where $S_{p}^{s,r}$ denotes the periodic dependency score for different object entities, ${tanh}$ is the activation function, $\oplus$ denotes the connection operator, and $\mathbf{W}_{p}\in{R}^{d\times 2d},\mathbf{b}_{p}\in{R}^{d\times 2d}$ are the trainable parameters. We compute the initial similarity score between each entity $o_{i}\in E$ and the query $q$ using a linear layer with activation.
For non-periodic dependencies, we calculate the similarity scores $d_{E}(s,o)=\parallel s-o\parallel$ using the following Euclidean distance formula to obtain the non-periodic dependency score $S_{np}^{s,r}$:
$\mathbf{S}_{np}^{s,r}(o_{i})=tanh(\mathbf{W}_{np}{s}\oplus r+\mathbf{b}_{np})\mathbf{E}^{T}-\mathbf{Z}_{t}^{s,r}+d_{{E}}(s,o_{i}),$ (12)
Method ICEWS18 ICEWS14
MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
ComplEx(2016) 30.09 21.88 34.15 45.96 24.47 16.13 27.49 41.09
R-GCN(2018) 23.19 16.36 25.34 36.48 26.31 18.23 30.43 45.34
ConvE(2018) 36.67 28.51 39.80 50.69 40.73 33.20 43.92 54.35
RE-Net(2019) 42.93 36.19 45.47 55.80 45.71 38.42 49.06 59.12
xERTE(2020) 36.95 30.71 40.38 49.76 32.92 26.44 36.58 46.05
CyGNet(2021) 17.21 30.97 46.85 27.12 27.43 42.63 57.90 37.65
EvoKG(2022) 29.67 12.92 33.08 58.32 18.30 6.30 19.43 39.37
RPC(2023) 34.91 24.34 38.74 55.89 44.55 34.87 49.80 65.08
CENET(2023) 43.72 37.91 45.65 54.79 51.93 48.24 52.81 59.33
DHE-TKG(2024) 29.23 19.15 33.31 40.02 30.13 44.99
RLGNet(2024) 29.90 20.18 33.64 49.08 39.06 29.34 42.03 58.12
HisRES(2024) 37.69 26.46 42.75 59.70 50.48 39.57 56.65 71.09
DiffuTKG(2024) 36.72 25.73 57.81 48.51 36.41 72.75
DPCL-Diff 47.02 40.03 49.93 60.66 66.59 62.89 67.79 73.64
Table 1: TKG entity extrapolation results on four datasets. The time-filtered MRR, H@1, H@3, and H@10 metrics are multiplied by 100. The best results are in bold, and the previous best results are underlined.
The training objective of learning from periodic and non-periodic events is to minimize the loss $\mathcal{L}_{\mathrm{ce}}$:
$\mathcal{L}_{ce}=-\sum_{q}\log\{\frac{exp(S_{p}^{s,r}(o_{i}))}{\sum\limits exp(S_{p}^{s,r}(o_{j}))}+\frac{exp(S_{np}^{s,r}(o_{i}))}{\sum\limits exp(S_{np}^{s,r}(o_{j}))}\},$ (13)
where $o_{i}$ denotes the ground truth object entity for a given query $q$. The goal of $\mathcal{L}_{ce}$ is to distinguish the ground truth from other values by comparing each scalar in $S_{p}^{s,r}$ and $S_{np}^{s,r}$.
The model employs supervised contrastive learning to enhance its discriminative power by enforcing consistency within the same class and differentiation between different classes, controlled by the supervised contrastive loss $\mathcal{L}_{sup}$:
$\mathcal{L}_{sup}=\sum_{q\in I}\frac{-1}{\mid P(i)\mid}\sum_{p\in P(i)}\log\frac{\exp(z_{q}\cdot z_{p}/\tau)}{\sum\exp(z_{q}\cdot z_{a}/\tau)}.$ (14)
where $\tau$ is the temperature parameter set to 0.1 in experiments,$I$ is the set of positive samples,$P(i)$ is the set of queries with the same label in the small batch, $z_{q}$ denotes the embedding vector of query $q$, $z_{p}$ represents the embedding vector of queries sharing the same label as $q$, $z_{a}$ denotes the embedding vector of all non-$q$ queries in the small batch.
DPCL-Diff will choose the object with the highest probability as the final prediction $o_{i}$.
3.4 Training Details and Inference
DPCL-Diff minimizes the loss function throughout model training:
$\mathcal{L}_{\mathrm{dpcl}}=\alpha\cdot\mathcal{L}_{\mathrm{diff}}+(1-\alpha)\cdot(\mathcal{L}_{\mathrm{ce}}+\mathcal{L}_{\mathrm{sup}}),$ (15)
where $\alpha$ is the equilibrium coefficient.
In GNDiff, the inference process closely mirrors the generation process, so candidate entity probability is computed via reverse diffusion.
$\mathbf{P}_{diff}(o_{i}\mid q)=\prod_{t=1}^{T}p_{\theta}(x_{t-1}\mid x_{t}),$ (16)
where $o_{i}$ is the tail entity representation in the generated sample $x_{t}$, and $x_{t}$ represents the sample corresponding to the query $q$.
For DPCL the probability distribution of entities is calculated based on the predicted missing objects:
$\mathbf{P}_{dpcl}(o_{i}\mid q)=\frac{\exp(s(o_{i}\mid q))}{\sum_{o_{j}\in\mathcal{E}}\exp(s(o_{j}\mid q))},$ (17)
The final probability $\mathbf{P}(o_{i})$ is obtained by averaging the probabilities from the diffusion process and the DPCL prediction:
$\mathbf{P}(o_{i}\mid q)=\frac{1}{2}\left(\mathbf{P}_{diff}(o_{i}\mid q)+\mathbf{P}_{dpcl}(o_{i}\mid q)\right).$ (18)
4 Experiments
Method WIKI YAGO
MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
ComplEx(2016) 47.84 38.15 50.08 58.46 61.29 54.88 62.28 69.42
R-GCN(2018) 37.57 28.15 39.66 —— 41.3 32.56 44.44 ——
ConvE(2018) 47.57 38.76 50.1 63.74 62.32 56.19 63.97 70.44
RE-Net(2019) 51.97 48.01 52.07 53.91 65.16 63.29 65.63 68.08
xERTE(2020) 58.75 58.46 58.85 59.34 58.75 58.46 58.85 60.48
CyGNet(2021) 47.89 66.44 78.70 58.78 17.21 30.97 46.85 27.12
EvoKG(2022) 50.66 12.21 63.84 64.73 55.11 54.37 81.38 92.73
CENET(2023) 66.87 66.77 66.93 66.98 79.64 78.65 79.91 81.42
DHE-TKG(2024) 51.20 57.47 69.25 62.93 71.00 82.72
RLGNet(2024) 64.34 61.03 66.71 69.51 80.17 76.52 83.57 84.96
DPCL-Diff 68.44 68.41 68.45 68.55 84.45 84.23 84.65 84.75
Table 2: Experimental results of temporal link prediction on two public KGs.
4.1 Experimental Setup
Datasets and Evaluation Metrics. To evaluate the effectiveness of the proposed DPCL-Diff, we report the reasoning performance on four public datasets: ICEWS14 Trivedi et al. (2017), ICEWS18 Jin et al. (2019), WIKI Leblay and Chekol (2018), and YAGO Mahdisoltani et al. (2013). Table 3 shows more detailed information about the datasets. We divided all datasets into three groups, roughly partitioning them into train (80%), validation (10%), and test (10%) sets by timestamp. In our experiments, we chose the mean reversed rank (MRR) and the hit rate @1,3,10 Liang et al. (2023) as the metrics.
Dataset #Ent #Re #Train #Valid #Test
ICEWS18 23,033 256 373,018 45,995 49,545
ICEWS14 12,498 260 535,654 63,788 65,861
WIKI 12,554 24 539,286 67,538 63,110
YAGO 10,623 10 161,540 19,523 20,026
Table 3: Statistics of the datasets.
Baselines. DPCL-Diff is compared with both static and temporal KG reasoning models. Static KG reasoning models, which do not consider temporal information, include ComplEx Trouillon et al. (2016), R-GCN Schlichtkrull et al. (2018) and ConvE Dettmers et al. (2018). In contrast, temporal KG reasoning models are designed to process temporal information and capture evolving patterns in TKG, such as RE-Net Jin et al. (2019), xERTE Z et al. (2020), CyGNet Zhu et al. (2021), EvoKG Park et al. (2022), RPC Liang et al. (2023), CENET Xu et al. (2023), DHE-TKG Liu et al. (2024b), RLGNet Lv et al. (2024), and DiffuTKG Cai et al. (2024).
Implementation Details. For all datasets, the initial embedding dimension is configured as 200. We set the batch size to 64, the embedding dimension to 200, and the learning rate to 0.001, and we used the Adam optimizer. The training calendar element limit for L is 30, and the comparison learning calendar element limit for the second stage is 20. For baseline settings, we use their recommended configuration. All experiments were implemented on servers with RTX 3090 GPUs equipped with Intel Xeon Gold 6330 CPUs.
Method YAGO ICEWS18
MRR Hits@1 Hits@5 Hits@10 MRR Hits@1 Hits@5 Hits@10
w/o- GNDif 81.30 80.64 81.42 82.26 44.93 38.22 47.53 57.08
w/o- DPCL 84.21 84.19 84.23 84.34 46.51 39.92 48.92 59.67
DPCL-Diff 84.45 84.23 84.65 84.75 47.02 40.03 49.93 60.66
Table 4: DPCL-Diff ablation studies on YAGO and ICEWS18 datasets.
4.2 Main Results
We compared the experimental results in Table 1 and Table 2, which provides the MRR and Hits@1/3/10 metrics of DPCL-Diff on four datasets. DPCL-Diff outperforms 13 other comparison models in most cases, particularly excelling on ICEWS14, where it achieves a 29.54% improvement in Hits@1 over our baseline model CENET. This significant gain is mainly due to effective data generation techniques, enabling the model to handle unseen events more effectively. ICEWS14 has a high proportion of new events(about 30% Xu et al. (2023)) and contains the most significant volume of data among the four datasets. DPCL-Diff generates substantial-high-quality data through GNDiff, enhancing inference on these events and improving overall performance. / Moreover, while DiffuTKG was the first to introduce diffusion processes in TKG inference, DPCL-Diff further leverages a pre-trained language model (PLM) in the denoising process, capturing contextual information more effectively and aligning data generation closer to actual event distributions. This is particularly advantageous in scenarios with sparse interaction trajectories, leading to superior predictive accuracy and allowing DPCL-Diff to outperform DiffuTKG across key metrics, especially in datasets with a high proportion of new events./ For the YAGO and WIKI datasets, DPCL-Diff also outperforms other algorithms on most metrics, except for Hits@10 on YAGO, which is slightly lower than RLGNet but significantly higher in Hits@1. DPCL-Diff ability to distinguish similar periodic event entities improves prediction accuracy. Although YAGO and WIKI have a lower proportion of new events (about 10% Xu et al. (2023)) and a higher percentage of periodic events, which limits GNDiff performance, DPCL-Given that the GNDiff parameters used in this experiment are relatively low, with the batch size set to 64, these settings may influence the results. We believe the performance could be significantly improved if the batch size is increased to 512 or 1024.
4.3 Ablation Study
To investigate the effectiveness of the modules in DPCL-Diff, we verified the effectiveness of GNDiff and DPCL through ablation experiments. The ablation results are shown in Table 4. Here, "w/o-GN Diff" denotes the removal of the graph node diffusion model, and "w/o-DPCL" denotes the removal of the Poincaré mapping, which maps both positive and negative samples into the Euclidean space for contrast learning. These results illustrate the effectiveness of our proposed GNDiff and DPCL. The performance degradation is particularly noticeable after the removal of GNDiff. This demonstrates that generating a large number of high-quality data related to the predicted event through GNDiff significantly affects event prediction. DPCL leverages Poincaré space properties to distinguish similar periodic events better, enhancing inference accuracy.
Method YAGO
MRR Hits@1 Hits@5 Hits@10
Euc/Hyp 81.20 80.60 81.26 82.08
Hyp/Hyp 81.23 80.62 81.30 82.14
Euc/Euc 79.64 78.65 79.91 81.42
Hyp/Euc 81.30 80.64 81.42 82.26
Table 5: Experimental results of DPCL under different spatial mapping strategies
4.4 Dual Domain Mapping Strategy Analysis
In this study, we separately examined the effects of different spatial mapping strategies on the learning effect of DPCL. Table 4 shows the results of the experiments on the YAGO dataset without the GNDiff. Euc/Hyp denotes that periodic and non-periodic event entities are mapped to Euclidean and Poincaré spaces respectively for comparative learning, Hyp/ Euc denotes that the two types of entities are mapped to Poincaré and Euclidean space, Hyp/Hyp denotes that both types of entities are mapped to Poincaré space, and Euc / Euc denotes that the two types of entities are mapped to Euclidean space respectively. From the experimental results, it can be seen that the Hyp/Euc method is the most effective, while the Euc/Euc method is the least effective. This suggests that mapping periodic and non-periodic event entities to separate spaces can improve the modeling of comparative learning. Specifically, the geometric properties of the Poincaré space can effectively distinguish entities whose feature representations are too close to each other in the Euclidean space.
5 Conclusion
In this paper, we propose a temporal knowledge graph inference method based on a graph node diffusion model with dual-domain periodic contrastive learning. This method aims to alleviate the prediction difficulties caused by the limited interactions of new events on the timeline. We enhance new event reasoning by generating graph data through graph node diffusion, where noise is introduced and adjusted to simulate real-world event mechanisms, producing samples that better reflect actual distributions. Additionally, the performance of the temporal knowledge reasoning task is enhanced by mapping event entities to Poincaré space and Euclidean space through dual-domain periodic contrastive learning, effectively distinguishing between periodic and non-periodic events. Extensive experiments demonstrate the reasoning advantages and effectiveness of DPCL-Diff on public temporal knowledge graph datasets.
6 Limitations
This study does not incorporate adaptive embedding strategies to distinguish between periodic and non-periodic events. Consequently, our approach lacks the flexibility to dynamically adjust to different types of temporal knowledge graph data. This limitation could potentially impact the model’s effectiveness when applied to datasets with varying temporal characteristics.
Appendices
7 Extended Experiments
7.1 Dual Domain embedded visual analytics
In this section, we investigate the visualization of entities in DPCL with bi-domain embedding. As shown in Fig. 2, we visualize two sets of entity pairs—(Germany, Berlin) and (Germany, Munich)—with the same relation, mapping them to Euclidean and Poincaré space, respectively. In Euclidean space, due to the similar Euclidean distance between Berlin and Munich to Germany, it is difficult to distinguish between these two entities effectively. However, in Poincaré space, Germany is significantly farther from Munich, bringing Germany closer to Berlin and making it easier to derive the correct inference. By leveraging the spatial properties of Poincaré space, we can more effectively distinguish similar periodic event entities and improve the overall prediction accuracy of the model.
Figure 2: Illustration of 2D entity embedding learned by Euc (left) and Hyp (right) on YAGO.

### v2

@@
-DPCL-Diff: The Temporal Knowledge Graph Reasoning based on Graph Node Diffusion Model with Dual-Domain Periodic Contrastive Learning
+DPCL-Diff: Temporal Knowledge Graph Reasoning Based on Graph Node Diffusion Model with Dual-Domain Periodic Contrastive Learning
 Abstract
@@
 This study also investigates the combined effectiveness of GNDiff and DPCL in TKG tasks.
-DPCL-Diff: The Temporal Knowledge Graph Reasoning based on Graph Node Diffusion Model with Dual-Domain Periodic Contrastive Learning
-Yukun Cao1, Lisheng Wang2, Luobing Huang3thanks: 123The authors are affiliated with the School of Computer Science and Technology, Shanghai University of Electric Power (SUEP), Shanghai 201306, China thanks: 1First author (email: marilyn_cao@163.com)thanks: 2Corresponding author (email: 2207992575@mail.shiep.edu.cn)
-1 Introduction
-Knowledge graphs (KGs) store a large amount of information about the natural world and have shown great success in many downstream applications, such as natural language processing Hu et al.
-(2022); Wang et al.
-(2020), recommendation system Xuan et al.
-(2023); Yu et al.
-(2022), and information retrieval Zhou et al.
-(2022).
+Introduction
+Knowledge graphs (KGs) store a large amount of information about the natural world and have shown great success in many downstream applications, such as natural language processing (Hu et al.
+2022; Wang et al.
+2020), recommendation system (Xuan et al.
+2023; Yu et al.
+2022), and information retrieval (Zhou et al.
+2022).
 Traditional KGs typically only contain static snapshots of facts, integrating facts (also known as events) in the form of static relational triples $(s,r,o)$, where $s$ and $o$ represent the subject and object entities, respectively, and $r$ represents the relationship type.
-However, in the real world, knowledge evolves and constantly exhibits complex temporal dynamics Wang et al.
-(2017); Yin and Cui (2016), which has inspired the construction and application of Temporal Knowledge Graph (TKG).
+However, in the real world, knowledge evolves and constantly exhibits complex temporal dynamics (Wang et al.
+2017; Yin and Cui 2016), which has inspired the construction and application of Temporal Knowledge Graph (TKG).
 TKG extends the previous static relational triples $(s,r,o)$ to quaternions $(s,r,o,t)$ with timestamps.
 Thus, the TKG consists of multiple snapshots where facts in the same snapshot appear simultaneously.
-Existing studies have identified two main types of new events in TKG: periodic events that will recur Liang et al.
-(2023); Li et al.
-(2021), and events that have not occurred before but may occur in the future Xu et al.
-(2023).
-For these new events, TKG reasoning offers fresh perspectives and insights for many downstream applications, such as event prediction Zhou et al.
-(2023), political decision making Zhang and Zhou (2022), dialogue generation Hou et al.
-(2024), and text generation Li et al.
-(2022b).
+Existing studies have identified two main types of new events in TKG: periodic events that will recur (Liang et al.
+2023; Li et al.
+2021), and events that have not occurred before but may occur in the future (Xu et al.
+2023).
+For these new events, TKG reasoning offers fresh perspectives and insights for many downstream applications, such as event prediction (Zhou et al.
+2023), political decision making (Zhang and Zhou 2022), dialogue generation (Hou et al.
+2024), and text generation (Li et al.
+2022b).
 These applications have greatly stimulated intense interest in TKG.
-In this work, we focus on predicting new facts in future time, a task known as graph extrapolation Chen et al.
-(2023); Li et al.
-(2021).
+In this work, we focus on predicting new facts in future time, a task known as graph extrapolation (Chen et al.
+2023; Li et al.
+2021).
 Our goal is to predict the missing entities in the query $(s,p,?,t)$ for future timestamps $t$ that still need to be observed in the training set.
-Most current work models the structural and temporal characteristics of TKG to capture the specific relationships and temporal dependencies between different events for future event prediction.Many studies Liang et al.
-(2023); Xu et al.
-(2023); Li et al.
-(2021)can predict repetitive or periodic events by referring to the known historical events and distinguishing the different effects of periodic and non-periodic events on reasoning tasks through contrastive learning.
+Most current work models the structural and temporal characteristics of TKG to capture the specific relationships and temporal dependencies between different events for future event prediction.Many studies (Liang et al.
+2023; Xu et al.
+2023; Li et al.
+2021)can predict repetitive or periodic events by referring to the known historical events and distinguishing the different effects of periodic and non-periodic events on reasoning tasks through contrastive learning.
 However, in actual reasoning tasks, some periodic events share the same head entities and relations, differing only in their tail entities.
 This results in overly similar representations, making it difficult to distinguish them during reasoning.
-Additionally, In event-based TKG, new events that have never occurred account for about 40% Jäger (2018).
+Additionally, In event-based TKG, new events that have never occurred account for about 40% (Jäger 2018).
 Due to the sparse traces of temporal interactions throughout the timeline of these new events, it is impossible to use reasoning methods based on high-frequency periodic events, resulting in poor reasoning performance for these types of events.
-To address the above problem, we propose a diffusion model-based dual-domain periodic contrast learning for timing knowledge graph reasoning (DPCL-Diff), which employs a Graph Node Diffusion model (GNDiff) to reason about new events.
-GNDiff tackles the issue of sparse interaction traces for new events on the timeline by introducing noise into existing correlated sparse events of the new entities.
-This diffusion mechanism simulates the real-world occurrence of new events and is gradually refined during the denoising process, guided by pre-trained language models (PLMs) that have been trained on extensive corpora and possess a foundational understanding of background knowledge.
-Consequently, this approach generates high-quality data that closely aligns with the actual distribution.
-The diffusion model Ho et al.
-(2020); Song et al.
-(2020) is a cutting-edge generative framework with demonstrated powerful capabilities in both image Ramesh et al.
-(2022); Rombach et al.
-(2022); Saharia et al.
-(2022)and text generation Ou and Jian (2024); Gong et al.
-(2022); Strudel et al.
-(2022).
-This provides more correlation information for inferring new events and dramatically improves the ability to reason about new events in TKG.
-For event inference tasks containing a large amount of temporal interaction information, we propose a novel dual-domain periodical contrast learning (DPCL) method.
-The periodic and non-periodic event entities related to the current event are mapped into the Poincaré and Euclidean spaces for contrastive learning, respectively.
-The Poincaré space property is utilized to distinguish the similar periodic event entities better.
+To address the above issue, we propose a diffusion model-based dual-doma in periodic contrast learning (DPCL-Diff) for temporal knowledge graph reasoning (TKG), using a Graph Node Diffusion model (GNDiff) to infer new events.
+GNDiff solves the sparse interaction traces for new events by adding noise to correlated sparse events of new entities on the timeline.
+This diffusion mechanism simulates the real-world occurrence of new events, generating high-quality data that closely align with the actual distribution through gradual correction during the denoising process.
+It also captures the event graph structure, enriching semantic representations and improving the model’s ability to reflect real-world dynamics.
+The diffusion model (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020) is a state-of-the-art generative framework with proven strength in both image (Ramesh et al.
+2022; Rombach et al.
+2022; Saharia et al.
+2022) and text generation (Ou and Jian 2024; Gong et al.
+2022; Strudel et al.
+2022).
+This enhances new event inference and reasoning in TKGs.
+For tasks with large temporal interaction data, we propose a novel Dual-Domain Periodic Contrastive Learning (DPCL) method.
+Here, periodic and non-periodic event entities are mapped into Poincaré and Euclidean spaces for contrastive learning, respectively.
+The Poincaré space is used to better distinguish similar periodic entities, capturing entity relationships more effectively than Euclidean space, leading to more accurate link prediction (Balazevic, Allen, and Hospedales 2019; Zeb et al.
+2024).
 The main contributions of this thesis are summarized below:
   • We propose a TKG model called DPCL-Diff. DPCL-Diff can reason not only about periodic events but also about new events through a diffusion generation mechanism.
-• GNDiff introduces a novel approach in TKG.
-GNDiff leverages pre-trained language models (PLMs) to conduct non-Markov decision chain-based diffusion for new events, generating high-quality data and effectively addressing the challenge of sparse interaction traces.
-• We propose the DPCL method, where periodic and non-periodic event entities are mapped into the Poincaré space and Euclidean space to better distinguish similar periodic event entities.
+  • GNDiff is introduced in the field of TKG.
+GNDiff can perform non-Markov decision chain-based graph node diffusion for new events, which generates a large amount of high-quality graph data related to new events.
+  • We propose the DPCL method, where periodic and non-periodic event entities are mapped into Poincaré and Euclidean spaces, respectively, to more accurately capture the relationships between periodic event entities and distinguish similar ones.
   • We conducted experiments on four public datasets.
@@
 Figure 1: Overall architecture of DPCL-Diff. The first part is GNDiff, and the second part is DPCL.
-2 Related Work
-2.1 Diffusion Model
-Diffusion models Ho et al.
-(2020) are latent variable models designed for continuous data, achieving state-of-the-art sample quality in image and audio generation.
-To handle discrete data, prior work explored text diffusion models in discrete state spaces and defined diffusion processes for such data Li et al.
-(2024); Lezama et al.
-(2022); Austin et al.
-(2021).
+Related Work
+Diffusion Model
+Diffusion models (Ho, Jain, and Abbeel 2020) are latent variable models for continuous data, achieving state-of-the-art image and audio generation quality.
+For discrete data, prior work explored text diffusion in discrete spaces (Li et al.
+2024; Lezama et al.
+2022; Austin et al.
+2021).
 Li et al.
-(2022a) focuses on continuous diffusion modelling for text, enabling gradient-based methods to control generation by introducing continuous latent representations.
-Ou and Jian (2024) combines discrete diffusion models with denoising techniques, significantly improving text generation quality and diversity in unconditional and conditional tasks.
-2.2 Temporal Knowledge Graph Reasoning
-TKG reasoning can be approached through interpolation or extrapolation Chen et al.
-(2023); Li et al.
-(2021).
-Interpolation addresses missing events between $t_{0}$ to $t_{n}$,while extrapolation predicts events beyond $t_{n}$.
-Interpolation models like TeAST Li et al.
-(2023), Re-Temp Wang et al.
-(2023), and TempCaps Fu et al.
-(2022) infer missing relationships but cannot predict future events.
-Extrapolation methods focus on future event prediction.
-Know-Evolve Trivedi et al.
-(2017), the first to learn nonlinear evolutionary entity embeddings, struggles with long-term dependencies.
-CyGNet Zhu et al.
-(2021) uses replication generation to identify frequent repetitive events, while CENET Xu et al.
-(2023) builds on this with contrastive learning to identify highly correlated entities.
-TARGAN Xie et al.
-(2023) enhances long-term modeling using attentional mechanisms.
-Other models, including DynaQ Liu et al.
-(2024a), EvoGNN Pareja et al.
-(2020), and TempoR Wu et al.
-(2020) have improved prediction but still face challenges in jointly capturing temporal evolution, semantic dependencies, and scalability.
-3 Method
+(2022a) enables gradient-based control of text diffusion with continuous latent representations.
+Ou and Jian (2024) combines discrete diffusion and denoising to improve text generation quality and diversity in both unconditional and conditional tasks.
+Temporal Knowledge Graph Reasoning
+TKG reasoning involves interpolation and extrapolation (Chen et al.
+2023; Li et al.
+2021).
+Interpolation fills missing events within $t_{0}$ to $t_{n}$, while extrapolation predicts events beyond $t_{n}$.
+Interpolation models like TeAST (Li, Su, and Gao 2023), Re-Temp (Wang, Han, and Poon 2023), and TempCaps (Fu et al.
+2022) address missing relationships but cannot predict future events.
+Extrapolation methods focus on future events: Know-Evolve (Trivedi et al.
+2017) captures nonlinear entity evolution but struggles with long-term dependencies, while CyGNet (Zhu et al.
+2021) and CENET (Xu et al.
+2023) leverage repetitive patterns and contrastive learning.
+TARGAN (Xie et al.
+2023) improves long-term modeling with attention mechanisms.
+Models like DynaQ (Liu et al.
+2024a), EvoGNN (Pareja et al.
+2020), and TempoR (Wu et al.
+2020) enhance prediction but face challenges in capturing temporal evolution, semantic dependencies, and scalability.
+Method
 As shown in Fig.
-1, for novel events, the DPCL-Diff model leverages GNDiff to inject noise into the sparse correlated events of the new event entity.
+1, for new events, the DPCL-Diff model leverages GNDiff to inject noise into the sparse correlated events of the new event entity.
 This approach simulates the real-world mechanism of new events, generating numerous high-quality data samples.
-For other events, the DPCL method identifies highly correlated entities through comparative learning and better distinguishes similar periodic event entities using Poincaré space properties.
-In the following few sections, we describe our proposed approach in detail.
-3.1 Preliminaries
+For other events, the DPCL method identifies highly correlated entities through contrastive learning and more accurately captures the relationships between periodic event entities using the properties of Poincaré space, thereby distinguishing similar periodic entities.
+In the following sections, we describe our proposed approach in detail.
+Preliminaries
 TKG $G=\{G_{0},G_{1},...,G_{t},...,G_{T}\}$ is a sequence of KGs, each $G_{t}=\{\mathcal{E},\mathcal{R},\mathcal{T}\}$ contains facts that occur at timestamp $t$.
@@
 Each fact is a quadruple $(s,r,o,t)$ , where $s\in\mathcal{E}$ is the subject entity, $o\in\mathcal{E}$ is the object entity, and $r\in\mathcal{R}$ is the relation at timestamp $t$ between $s$ and $o$ that occurs, the dimension is $d$.
-||×$\mathcal{E}\in\mathbb{R}^{|\mathcal{E}|\times d}$ is the embedding of all entities, and its rows represent the embedding vectors of an entity, such as $s$ and $o$.
-Similarly,||×$\mathcal{R}\in\mathbb{R}^{|\mathcal{E}|\times d}$ is the embedding of all relation types.
+$\textbf{E}\in\mathbb{R}^{|\mathcal{E}|\times d}$ is the embedding of all entities, and its rows represent the embedding vectors of an entity, such as $s$ and $o$.
+Similarly, $\textbf{R}\in\mathbb{R}^{|\mathcal{R}|\times d}$ is the embedding of all relation types.
 The TKG inference task aims to predict missing objects by answering a query $q=(s,r,?,t)$ using a given history of TKG.
@@
 The following sections detail these modules.
-3.2 Graph Node Diffusion Model
-In this section, we introduce GNDiff, showcasing the potential of diffusion models for TKG applications.
-While previous work has effectively explored the diffusion of discrete textual data, applying diffusion models to the discrete graph nodes in TKGs can be similarly effective.
-Traditional reasoning methods, which rely on historical data to learn implicit associations, need help to reason about new events due to the sparse historical data associated with them.
-In contrast, GNDiff employs a diffusion model to simulate the event generation mechanism, producing many high-quality samples for new event inference.
-GNDiff consists of a forward diffusion process and a reverse diffusion process.
-For each quadruple $(s,r,o,t)$, after obtaining the vector representations of $s$, $r$, and $o$, these vectors are concatenated into a single vector, which serves as the input to the graph node diffusion model.
-We then obtain the sample data $x_{0}\sim q(x_{0})$.
-Since TKGs involve discrete data, each element of $x_{t}$ is a discrete random variable with $K$ categories.
-For graph node, $K=\left|\mathcal{V}\right|$ represents the number of node types.
-Representing $x_{t}$ as a stack of heat vectors, the process of adding noise is denoted as shown in Equation 3:
+Graph Node Diffusion Model
+In this section, we introduce GNDiff, demonstrating the effectiveness of diffusion models in TKGs.
+While diffusion models have been applied to discrete textual data, they are equally effective for discrete graph nodes in TKGs.
+Traditional methods, reliant on historical data, struggle with new events due to sparse data.
+In contrast, GNDiff uses a diffusion model to simulate event generation, producing high-quality samples for new event inference.
+GNDiff includes a forward and a reverse diffusion process.
+For each quadruple $(s,r,o,t)$, the vector representations of $s$, $r$, and $o$ are concatenated, forming a single vector as input to the graph node diffusion model.
+The sample data $x_{0}\sim q(x_{0})$ is then obtained.
+Since TKGs are discrete, each element of $x_{t}$ is a discrete random variable with $K$ categories, where $K=\left|\mathcal{V}\right|$ is the number of node types.
+The noise addition process is represented by a stack of heat vectors as shown in Equation 3:
 $q(x_{t}\mid x_{t-1})=\mathrm{Cat}(x_{t};p=x_{t-1}Q_{t}),$ (3)
 where $\mathrm{Cat}(\cdot)$ denotes that the noise addition process is based on the category distribution, and ${Q}_{t}$ is used to control the transition matrix of the node from one state to another:
-$[Q_{t}]_{i,j}=q(x_{t}=j\mid x_{t-1}=i),$ (4)
+$[Q_{t}]_{i,j}=q(x_{t}=j\mid x_{t-1}=i).$ (4)
 The conversion relationship is shown in Equation 5:
@@
 Note that $\odot$ denotes elementwise multiplication, and division is performed rowwise.
-To be combined with pre-trained denoising language models, we incorporate an absorbing state, allowing for the generation of samples during denoising.
-We use the masking mechanism to simulate the possible inactive state of the node.
+To incorporate the pre-trained language model into the denoising process, we introduce a specific ”inactive” state as the absorbing state, allowing for the generation of samples during denoising.
+We use the mask mechanism to simulate the possible inactive state of the node.
 The transition matrix adjustment process is shown in Equation 6:
-$[Q_{t}]_{i,j}=\begin{cases}1&\mathrm{if}\quad i=j=[\mathrm{M}],\\ \beta_{t}&\mathrm{if}\quad j=[\mathrm{M}],\ \mathrm{i}\neq[\mathrm{M}],\\ 1-\beta_{t}&\mathrm{if}\quad i=j\neq[\mathrm{M}],\end{cases}$ (6)
+$[Q_{t}]_{i,j}=\begin{cases}1&\mathrm{if}\quad i=j=[\mathrm{M}],\\ \beta_{t}&\mathrm{if}\quad j=[\mathrm{M}],\ \mathrm{i}\neq[\mathrm{M}],\\ 1-\beta_{t}&\mathrm{if}\quad i=j\neq[\mathrm{M}].\end{cases}$ (6)
 Here, $[\mathrm{M}]$ denotes a node in a masked state (an inactive node), allowing nodes to transition naturally in and out of this state as the graph evolves, eventually converging to a smooth distribution $q(x_{T})$.
@@
 Samples are then generated through a backward diffusion process:
-$p_{\theta}({x}_{0:T})=p({x}_{T})\prod_{t=1}^{T}p_{\theta}({x}_{t-1}\mid{x}_{t},t),$ (7)
+$p_{\theta}({x}_{0:T})=p({x}_{T})\prod\nolimits_{t=1}^{T}p_{\theta}({x}_{t-1}\mid{x}_{t},t).$ (7)
 We optimize the inverse process by minimizing the variational lower bound to maximize the log-likelihood of the original data.
@@
 $\mathcal{L}_{\mathrm{diff}}=\mathbb{E}_{q}[\mathrm{D}_{\mathrm{KL}}(q(x_{T}\mid x_{0})\parallel p_{\theta}(x_{T}))]+$ (8)
-$\mathbb{E}_{q}[\sum_{t=2}^{T}\mathrm{D}_{\mathrm{KL}}(q({x}_{t-1}|{x}_{t},{x}_{0})\parallel p_{\theta}({x}_{t-1}|{x}_{t},t))]$
+$\mathbb{E}_{q}\left[\sum\nolimits_{t=2}^{T}\mathrm{D}_{\mathrm{KL}}(q({x}_{t-1}\mid{x}_{t},{x}_{0})\parallel p_{\theta}({x}_{t-1}\mid{x}_{t},t))\right]$
 $-\log p_{\theta}(x_{0}\mid x_{1}),$
 where ${E}_{q}$ denotes the expectation under a noisy data distribution. $\mathrm{D}_{\mathrm{KL}}$ stands for Kullback-Leibler divergence.
-3.3 Dual-Domain Periodic Contrastive Learning
+To better control noise addition, GNDiff employs a noise schedule (He et al.
+2022) that measures noise at each step by masking and evenly distributing information.
+Tokens are assigned varying mask probabilities, ordered by informativeness into $T$ stages.
+The forward process begins by masking the most informative tokens and ends with the least informative, simplifying the reverse process.
+Specifically, mask information is evenly distributed during the forward step, as shown in Equation 9:
+$\sum\nolimits_{i=1}^{n}\overline{\alpha}_{t}^{i}H(x_{0}^{i})=\left(1-\frac{t}{T}\right)\sum\nolimits_{i=1}^{n}H(x_{0}^{i}),$ (9)
+where $H$ represents entropy, quantifying the information content of a node.
+$x_{0}^{i}$ denotes the $i$th token in the sequence and $n$ is the sequence length.
+$\overline{\alpha}_{t}^{i}=\prod_{j=1}^{t}(1-\beta_{j}^{i})$ is the probability that the node retains its original state at time step $t$, i.e.
+$x_{t}^{i}=x_{0}^{i}$.
+We expect that if $H(x_{t}^{i})<H(x_{t}^{j})$, then $\bar{\alpha}_{t}^{i}>\bar{\alpha}_{t}^{j}$, so low-information tokens appear earlier in the reverse process than high-information ones.
+In practice, a token’s entropy $H(x)$ is computed as the negative logarithm of its frequency in the training corpus.
+Based on these properties, we construct $\overline{\alpha}_{t}^{i}$ as follows:
+$\overline{\alpha}_{t}^{i}=1-\frac{t}{T}-G(t)\cdot\tilde{H}({x}_{0}^{i}),$ (10)
+$G(t)=\mu\sin\frac{t\pi}{T},$ (11)
+$\tilde{H}({x}_{0}^{i})=1-\frac{\sum_{j=1}^{n}H({x}_{0}^{j})}{nH({x}_{0}^{i})},$ (12)
+where $G(t)$ controls the influence of information at time step $t$, with its effect modulated by the hyperparameter $\mu$.
+Dual-Domain Periodic Contrastive Learning
 Inspired by Xu et al.
 (2023), we propose a dual-domain periodic contrastive learning method.
-Unlike the original approach, we use a dual-domain mapping strategy because we observed that specific queries involve multiple similar periodic events that are challenging to distinguish.
-To address this issue, we map entities from periodic and non-periodic events to Poincaré and Euclidean space, respectively, to identify highly correlated entities within each domain.
-By leveraging the spatial characteristics of Poincaré space, we effectively differentiate similar periodic entities.
+Unlike the original approach, we use a dual-domain mapping strategy because we observed that specific queries involve multiple periodic events that are challenging to distinguish.
+To address this issue, we map entities from periodic and non-periodic events to Poincaré and Euclidean spaces, respectively, to identify highly correlated entities within each domain.
+By leveraging the spatial characteristics of Poincaré space, we more accurately capture the relationships between periodic entities and distinguish similar ones.
 During data preprocessing, we analyze entity frequencies for a given query $q=(s,r,?,t)$, as shown in Equation 13:
 $\mathbf{Z}_{t}^{s,r}(o)=\lambda\cdot(\Phi_{\beta>0}-\Phi_{\beta=0}),$ (13)
-$\beta=\sum_{(s,r,o_{i},k)\in G_{k}}(o_{i}=o),$ (10)
+$\beta=\sum\nolimits_{(s,r,o_{i},k)\in G_{k}}(o_{i}=o),$ (14)
 where the value of each entity frequency is constrained by the hyperparameter $\lambda$.
@@
 The formula for calculating the Poincaré distance is provided as follows:
+$\mathbf{d}_{p}(s,o)=cosh^{-1}\Bigg{(}1+2\frac{\parallel s-o\parallel_{2}^{2}}{\left(1-\parallel s\parallel_{2}^{2}\right)\left(1-\parallel o\parallel_{2}^{2}\right)}\Bigg{)},$ (15)
 $\mathbf{S}_{p}^{s,r}(o_{i})=tanh(\mathbf{W}_{p}{s}\oplus r+\mathbf{b}_{p})\mathbf{E}^{T}+\mathbf{Z}_{t}^{s,r}+\mathbf{d}_{p}(s,o_{i}),$ (16)
@@
 For non-periodic dependencies, we calculate the similarity scores $d_{E}(s,o)=\parallel s-o\parallel$ using the following Euclidean distance formula to obtain the non-periodic dependency score $S_{np}^{s,r}$:
-$\mathbf{S}_{np}^{s,r}(o_{i})=tanh(\mathbf{W}_{np}{s}\oplus r+\mathbf{b}_{np})\mathbf{E}^{T}-\mathbf{Z}_{t}^{s,r}+d_{{E}}(s,o_{i}),$ (12)
-Method ICEWS18 ICEWS14
-MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
-ComplEx(2016) 30.09 21.88 34.15 45.96 24.47 16.13 27.49 41.09
-R-GCN(2018) 23.19 16.36 25.34 36.48 26.31 18.23 30.43 45.34
-ConvE(2018) 36.67 28.51 39.80 50.69 40.73 33.20 43.92 54.35
-RE-Net(2019) 42.93 36.19 45.47 55.80 45.71 38.42 49.06 59.12
-xERTE(2020) 36.95 30.71 40.38 49.76 32.92 26.44 36.58 46.05
-CyGNet(2021) 17.21 30.97 46.85 27.12 27.43 42.63 57.90 37.65
-EvoKG(2022) 29.67 12.92 33.08 58.32 18.30 6.30 19.43 39.37
+$\mathbf{S}_{np}^{s,r}(o_{i})=tanh(\mathbf{W}_{np}{s}\oplus r+\mathbf{b}_{np})\mathbf{E}^{T}-\mathbf{Z}_{t}^{s,r}+d_{{E}}(s,o_{i}).$ (17)
+The training objective of learning from periodic and non-periodic events is to minimize the loss $\mathcal{L}_{\mathrm{ce}}$:
+$\mathcal{L}_{\mathrm{ce}}=-\sum_{q\in I}\log\{\frac{exp(S_{p}^{s,r}(o_{i}))}{\sum\limits_{o_{j}\in\mathcal{E}}exp(S_{p}^{s,r}(o_{j}))}+\frac{exp(S_{np}^{s,r}(o_{i}))}{\sum\limits_{o_{j}\in\mathcal{E}}exp(S_{np}^{s,r}(o_{j}))}\}.$ (18)
+Here, $o_{i}$ denotes the ground truth object entity for a given query $q$.
+The goal of $\mathcal{L}_{ce}$ is to distinguish the ground truth from other values by comparing each scalar in $S_{p}^{s,r}$ and $S_{np}^{s,r}$.
+The model employs supervised contrastive learning to enhance its discriminative power by enforcing consistency within the same class and differentiation between different classes, controlled by the supervised contrastive loss $\mathcal{L}_{sup}$:
+$\mathcal{L}_{sup}=\sum_{q\in I}\frac{-1}{\mid P(i)\mid}\sum_{p\in P(i)}\log\frac{\exp(z_{q}\cdot z_{p}/\tau)}{\sum\exp(z_{q}\cdot z_{a}/\tau)},$ (19)
+where $I$ is the set of positive samples, $P(i)$ is the set of queries with the same label in the small batch, $z_{q}$ denotes the embedding vector of query $q$, $z_{p}$ represents the embedding vector of queries sharing the same label as $q$, $z_{a}$ denotes the embedding vector of all non-$q$ queries in the small batch.
+Method ICEWS18 ICEWS14 WIKI
+MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
+ComplEx(2016) 30.09 21.88 34.15 45.96 24.47 16.13 27.49 41.09 47.84 38.15 50.08 58.46
+R-GCN(2018) 23.19 16.36 25.34 36.48 26.31 18.23 30.43 45.34 37.57 28.15 39.66 ——
+ConvE(2018) 36.67 28.51 39.80 50.69 40.73 33.20 43.92 54.35 47.57 38.76 50.1 63.74
+RE-Net(2019) 42.93 36.19 45.47 55.80 45.71 38.42 49.06 59.12 51.97 48.01 52.07 53.91
+xERTE(2020) 36.95 30.71 40.38 49.76 32.92 26.44 36.58 46.05 58.75 58.46 58.85 59.34
+CyGNet(2021) 46.69 40.58 49.82 57.14 48.63 41.77 52.50 60.29 45.5 50.48 50.79 68.95
+EvoKG(2022) 29.67 12.92 33.08 58.32 18.30 6.30 19.43 39.37 50.66 12.21 63.84 64.73
 RPC(2023) 34.91 24.34 38.74 55.89 44.55 34.87 49.80 65.08
-CENET(2023) 43.72 37.91 45.65 54.79 51.93 48.24 52.81 59.33
-DHE-TKG(2024) 29.23 19.15 33.31 40.02 30.13 44.99
-RLGNet(2024) 29.90 20.18 33.64 49.08 39.06 29.34 42.03 58.12
+CENET(2023) 43.72 37.91 45.65 54.79 51.93 48.24 52.81 59.33 66.87 66.77 66.93 66.98
+DHE-TKG(2024) 29.23 19.15 33.31 40.02 30.13 44.99 51.20 57.47 69.25
+RLGNet(2024) 29.90 20.18 33.64 49.08 39.06 29.34 42.03 58.12 64.34 61.03 66.71 69.51
 HisRES(2024) 37.69 26.46 42.75 59.70 50.48 39.57 56.65 71.09
-DiffuTKG(2024) 36.72 25.73 57.81 48.51 36.41 72.75
-DPCL-Diff 47.02 40.03 49.93 60.66 66.59 62.89 67.79 73.64
-Table 1: TKG entity extrapolation results on four datasets.
+DPCL-Diff 47.02 40.03 49.93 60.66 66.59 62.89 67.79 73.64 68.44 68.41 68.45 68.55
+Improve(%) 7.55 $\uparrow$ 5.59 $\uparrow$ 9.38$\uparrow$ 10.71$\uparrow$ 28.23 $\uparrow$ 30.37 $\uparrow$ 28.37 $\uparrow$ 24.12 $\uparrow$ 2.35 $\uparrow$ 2.46 $\uparrow$ 2.26 $\uparrow$ 2.34 $\uparrow$
+Table 1: Experimental results of temporal link prediction on three datasets TKGs.
 The time-filtered MRR, H@1, H@3, and H@10 metrics are multiplied by 100.
 The best results are in bold, and the previous best results are underlined.
-The training objective of learning from periodic and non-periodic events is to minimize the loss $\mathcal{L}_{\mathrm{ce}}$:
-$\mathcal{L}_{ce}=-\sum_{q}\log\{\frac{exp(S_{p}^{s,r}(o_{i}))}{\sum\limits exp(S_{p}^{s,r}(o_{j}))}+\frac{exp(S_{np}^{s,r}(o_{i}))}{\sum\limits exp(S_{np}^{s,r}(o_{j}))}\},$ (13)
-where $o_{i}$ denotes the ground truth object entity for a given query $q$.
-The goal of $\mathcal{L}_{ce}$ is to distinguish the ground truth from other values by comparing each scalar in $S_{p}^{s,r}$ and $S_{np}^{s,r}$.
-The model employs supervised contrastive learning to enhance its discriminative power by enforcing consistency within the same class and differentiation between different classes, controlled by the supervised contrastive loss $\mathcal{L}_{sup}$:
-$\mathcal{L}_{sup}=\sum_{q\in I}\frac{-1}{\mid P(i)\mid}\sum_{p\in P(i)}\log\frac{\exp(z_{q}\cdot z_{p}/\tau)}{\sum\exp(z_{q}\cdot z_{a}/\tau)}.$ (14)
-where $\tau$ is the temperature parameter set to 0.1 in experiments,$I$ is the set of positive samples,$P(i)$ is the set of queries with the same label in the small batch, $z_{q}$ denotes the embedding vector of query $q$, $z_{p}$ represents the embedding vector of queries sharing the same label as $q$, $z_{a}$ denotes the embedding vector of all non-$q$ queries in the small batch.
-DPCL-Diff will choose the object with the highest probability as the final prediction $o_{i}$.
-3.4 Training Details and Inference
+Improve refers to the improvement of DPCL-Diff compared to our baseline CENET.
+Training Details and Inference
 DPCL-Diff minimizes the loss function throughout model training:
-$\mathcal{L}_{\mathrm{dpcl}}=\alpha\cdot\mathcal{L}_{\mathrm{diff}}+(1-\alpha)\cdot(\mathcal{L}_{\mathrm{ce}}+\mathcal{L}_{\mathrm{sup}}),$ (15)
+$\mathcal{L}_{dpcl}=\alpha\cdot\mathcal{L}_{diff}+(1-\alpha)\cdot(\mathcal{L}_{ce}+\mathcal{L}_{sup}),$ (20)
 where $\alpha$ is the equilibrium coefficient.
 In GNDiff, the inference process closely mirrors the generation process, so candidate entity probability is computed via reverse diffusion.
-$\mathbf{P}_{diff}(o_{i}\mid q)=\prod_{t=1}^{T}p_{\theta}(x_{t-1}\mid x_{t}),$ (16)
+$\mathbf{P}_{\text{diff}}(o_{i}\mid q)=\prod\nolimits_{t=1}^{T}p_{\theta}(x_{t-1}\mid x_{t}),$ (21)
 where $o_{i}$ is the tail entity representation in the generated sample $x_{t}$, and $x_{t}$ represents the sample corresponding to the query $q$.
 For DPCL the probability distribution of entities is calculated based on the predicted missing objects:
-$\mathbf{P}_{dpcl}(o_{i}\mid q)=\frac{\exp(s(o_{i}\mid q))}{\sum_{o_{j}\in\mathcal{E}}\exp(s(o_{j}\mid q))},$ (17)
+$\mathbf{P}_{dpcl}(o_{i}\mid q)=\frac{\exp(s(o_{i}\mid q))}{\sum_{o_{j}\in\mathcal{E}}\exp(s(o_{j}\mid q))}.$ (22)
 The final probability $\mathbf{P}(o_{i})$ is obtained by averaging the probabilities from the diffusion process and the DPCL prediction:
 $\mathbf{P}(o_{i}\mid q)=\frac{1}{2}\left(\mathbf{P}_{diff}(o_{i}\mid q)+\mathbf{P}_{dpcl}(o_{i}\mid q)\right).$ (23)
-4 Experiments
-Method WIKI YAGO
-MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
-ComplEx(2016) 47.84 38.15 50.08 58.46 61.29 54.88 62.28 69.42
-R-GCN(2018) 37.57 28.15 39.66 —— 41.3 32.56 44.44 ——
-ConvE(2018) 47.57 38.76 50.1 63.74 62.32 56.19 63.97 70.44
-RE-Net(2019) 51.97 48.01 52.07 53.91 65.16 63.29 65.63 68.08
-xERTE(2020) 58.75 58.46 58.85 59.34 58.75 58.46 58.85 60.48
-CyGNet(2021) 47.89 66.44 78.70 58.78 17.21 30.97 46.85 27.12
-EvoKG(2022) 50.66 12.21 63.84 64.73 55.11 54.37 81.38 92.73
-CENET(2023) 66.87 66.77 66.93 66.98 79.64 78.65 79.91 81.42
-DHE-TKG(2024) 51.20 57.47 69.25 62.93 71.00 82.72
-RLGNet(2024) 64.34 61.03 66.71 69.51 80.17 76.52 83.57 84.96
-DPCL-Diff 68.44 68.41 68.45 68.55 84.45 84.23 84.65 84.75
-Table 2: Experimental results of temporal link prediction on two public KGs.
-4.1 Experimental Setup
+DPCL-Diff will choose the object with the highest probability as the final prediction $o_{i}$.
+Experiments
+Experimental Setup
 Datasets and Evaluation Metrics.
-To evaluate the effectiveness of the proposed DPCL-Diff, we report the reasoning performance on four public datasets: ICEWS14 Trivedi et al.
-(2017), ICEWS18 Jin et al.
-(2019), WIKI Leblay and Chekol (2018), and YAGO Mahdisoltani et al.
-(2013).
+To evaluate the effectiveness of the proposed DPCL-Diff, we report the reasoning performance on four public datasets: ICEWS14 (Trivedi et al.
+2017), ICEWS18 (Jin et al.
+2019), WIKI (Leblay and Chekol 2018), and YAGO (Mahdisoltani, Biega, and Suchanek 2013).
+We divided all datasets into three groups, roughly partitioning them into train (80%), validation (10%), and test (10%) sets by timestamp.
+The partitioning of the ICEWS14 dataset differs slightly from the original dataset.
 Table 2 shows more detailed information about the datasets.
-We divided all datasets into three groups, roughly partitioning them into train (80%), validation (10%), and test (10%) sets by timestamp.
-In our experiments, we chose the mean reversed rank (MRR) and the hit rate @1,3,10 Liang et al.
-(2023) as the metrics.
+In our experiments, we chose the mean reversed rank (MRR) and the hit rate @1,3,10 (Liang et al.
+2023) as the metrics.
 Dataset #Ent #Re #Train #Valid #Test
@@
 Table 2: Statistics of the datasets.
+Method YAGO
+MRR Hits@1 Hits@3 Hits@10
+ComplEx(2016) 61.29 54.88 62.28 69.42
+R-GCN(2018) 41.3 32.56 44.44 ——
+ConvE(2018) 62.32 56.19 63.97 70.44
+RE-Net(2019) 65.16 63.29 65.63 68.08
+xERTE(2020) 58.75 58.46 58.85 60.48
+CyGNet(2021) 63.47 64.26 65.71 68.95
+EvoKG(2022) 55.11 54.37 81.38 92.73
+CENET(2023) 79.64 78.65 79.91 81.42
+DHE-TKG(2024) 62.93 71.00 82.72
+RLGNet(2024) 80.17 76.52 83.57 84.96
+DPCL-Diff 84.45 84.23 84.65 84.75
+Improve(%) 6.04 $\uparrow$ 7.09 $\uparrow$ 5.86 $\uparrow$ 4.09 $\uparrow$
+Table 3: Experimental results of temporal link prediction on YAGO.
 Baselines.
 DPCL-Diff is compared with both static and temporal KG reasoning models.
-Static KG reasoning models, which do not consider temporal information, include ComplEx Trouillon et al.
-(2016), R-GCN Schlichtkrull et al.
-(2018) and ConvE Dettmers et al.
-(2018).
-In contrast, temporal KG reasoning models are designed to process temporal information and capture evolving patterns in TKG, such as RE-Net Jin et al.
-(2019), xERTE Z et al.
-(2020), CyGNet Zhu et al.
-(2021), EvoKG Park et al.
-(2022), RPC Liang et al.
-(2023), CENET Xu et al.
-(2023), DHE-TKG Liu et al.
-(2024b), RLGNet Lv et al.
-(2024), and DiffuTKG Cai et al.
-(2024).
+Static KG reasoning models, which do not consider temporal information, include ComplEx (Trouillon et al.
+2016), R-GCN (Schlichtkrull et al.
+2018) and ConvE (Dettmers et al.
+2018).
+In contrast, temporal KG reasoning models are designed to process temporal information and capture evolving patterns in TKG, such as RE-Net (Jin et al.
+2019), xERTE (Z et al.
+2020), CyGNet (Zhu et al.
+2021), EvoKG (Park et al.
+2022), RPC (Liang et al.
+2023), CENET (Xu et al.
+2023), DHE-TKG (Liu et al.
+2024b), RLGNet (Lv et al.
+2024), and HisRES (Zhang et al.
+2024).
+Method ICEWS18 YAGO
+MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
+w/o- GNDif 44.93 38.22 47.53 57.08 81.30 80.64 81.42 82.26
+w/o- DPCL 46.51 39.92 48.92 59.67 84.21 84.19 84.23 84.34
+DPCL-Diff 47.02 40.03 49.93 60.66 84.45 84.23 84.65 84.75
+Table 4: DPCL-Diff ablation studies on YAGO and ICEWS18 datasets.
+Method ICEWS14(new events) ICEWS18(new events) WIKI(new events)
+MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
+GNDiff 47.36 41.56 49.04 58.72 37.46 31.64 39.71 48.42 62.51 61.97 62.77 63.37
+DPCL 42.37 36.41 44.78 53.03 33.93 29.47 35.21 42.18 57.16 56.81 57.28 57.58
+Table 5: Performance of Graph Diffusion Models on New Events in ICEWS14 and ICEWS18 and WIKI.
+Method YAGO(new events)
+MRR Hits@1 Hits@3 Hits@10
+GNDiff 79.49 79.44 79.48 79.54
+DPCL 71.88 71.78 71.89 71.98
+Table 6: Performance of Graph Diffusion Models on New Events in YAGO.
+New Events Dataset #Train #Valid #Test
+ICEWS18 221,000 31,407 33,211
+ICEWS14 275,762 35,744 34,641
+WIKI 113,629 30,317 28,811
+YAGO 44,949 10,704 10,985
+Table 7: Statistics of New Events Datasets.
 Implementation Details.
-For all datasets, the initial embedding dimension is configured as 200.
-We set the batch size to 64, the embedding dimension to 200, and the learning rate to 0.001, and we used the Adam optimizer.
+For all datasets, DPCL uses an embedding dimension of 200, while GNDiff uses 128.
+We set the batch size to 64, the learning rate to 0.001, and used the Adam optimizer.
 The training calendar element limit for $L$ is 30, and the comparison learning calendar element limit for the second stage is 20.
+The value of the hyperparameter $\alpha$ is set to 0.2, and $\lambda$ is set to 2.
+Additionally, $\tau$ is the temperature parameter set to 0.1 in the supervised contrastive loss $\mathcal{L}_{sup}$.
 For baseline settings, we use their recommended configuration.
 All experiments were implemented on servers with RTX 3090 GPUs equipped with Intel Xeon Gold 6330 CPUs.
-Method YAGO ICEWS18
-MRR Hits@1 Hits@5 Hits@10 MRR Hits@1 Hits@5 Hits@10
-w/o- GNDif 81.30 80.64 81.42 82.26 44.93 38.22 47.53 57.08
-w/o- DPCL 84.21 84.19 84.23 84.34 46.51 39.92 48.92 59.67
-DPCL-Diff 84.45 84.23 84.65 84.75 47.02 40.03 49.93 60.66
-Table 4: DPCL-Diff ablation studies on YAGO and ICEWS18 datasets.
-4.2 Main Results
-We compared the experimental results in Table 1 and Table 2, which provides the MRR and Hits@1/3/10 metrics of DPCL-Diff on four datasets.
-DPCL-Diff outperforms 13 other comparison models in most cases, particularly excelling on ICEWS14, where it achieves a 29.54% improvement in Hits@1 over our baseline model CENET.
-This significant gain is mainly due to effective data generation techniques, enabling the model to handle unseen events more effectively.
-ICEWS14 has a high proportion of new events(about 30% Xu et al.
-(2023)) and contains the most significant volume of data among the four datasets.
-DPCL-Diff generates substantial-high-quality data through GNDiff, enhancing inference on these events and improving overall performance.
-/ Moreover, while DiffuTKG was the first to introduce diffusion processes in TKG inference, DPCL-Diff further leverages a pre-trained language model (PLM) in the denoising process, capturing contextual information more effectively and aligning data generation closer to actual event distributions.
-This is particularly advantageous in scenarios with sparse interaction trajectories, leading to superior predictive accuracy and allowing DPCL-Diff to outperform DiffuTKG across key metrics, especially in datasets with a high proportion of new events./ For the YAGO and WIKI datasets, DPCL-Diff also outperforms other algorithms on most metrics, except for Hits@10 on YAGO, which is slightly lower than RLGNet but significantly higher in Hits@1.
-DPCL-Diff ability to distinguish similar periodic event entities improves prediction accuracy.
-Although YAGO and WIKI have a lower proportion of new events (about 10% Xu et al.
-(2023)) and a higher percentage of periodic events, which limits GNDiff performance, DPCL-Given that the GNDiff parameters used in this experiment are relatively low, with the batch size set to 64, these settings may influence the results.
-We believe the performance could be significantly improved if the batch size is increased to 512 or 1024.
-4.3 Ablation Study
+Main Results
+We compared the experimental results in Table 1 and Table 3, presenting the MRR and Hits@1/3/10 metrics of DPCL-Diff across four datasets.
+DPCL-Diff consistently outperforms 12 baseline models, with a notable 30.37%(Hits@1) improvement over CENET on ICEWS14.
+This substantial gain is not solely attributable to the effective data generation of GNDiff for handling unseen events but also to its ability to generalize on periodic events.
+ICEWS14 high proportion of new events and large data volume amplify GNDiff impact, while the pre-trained language model (PLM) further enhances contextual representation.
+Additionally, DPCL-Diff’s ability to model structural and semantic relationships across events improves performance on both new and periodic events, with the sensitivity of the Hits@1 metric to ranking adjustments further emphasizing the improvement.
+On the YAGO and WIKI datasets, DPCL-Diff also achieves strong overall performance, with significantly higher Hits@1 compared to other models, even though Hits@10 on YAGO is slightly lower than RLGNet.
+The model’s ability to capture relationships between periodic events enhances prediction accuracy despite the lower proportion of new events and a higher proportion of periodic events, which limit GNDiff’s direct impact.
+Furthermore, the experiments were conducted with a relatively small batch size of 64, potentially constraining the model’s performance.
+Increasing the batch size to 512 or 1024 could further enhance its capacity to capture contextual and relational information, especially in datasets with more complex event patterns.
+Ablation Study
 To investigate the effectiveness of the modules in DPCL-Diff, we verified the effectiveness of GNDiff and DPCL through ablation experiments.
 The ablation results are shown in Table 4.
-Here, "w/o-GN Diff" denotes the removal of the graph node diffusion model, and "w/o-DPCL" denotes the removal of the Poincaré mapping, which maps both positive and negative samples into the Euclidean space for contrast learning.
+Here, ”w/o-GN Diff” denotes the removal of the graph node diffusion model, and ”w/o-DPCL” denotes the removal of the Poincaré mapping, which maps both positive and negative samples into the Euclidean space for contrast learning.
 These results illustrate the effectiveness of our proposed GNDiff and DPCL.
@@
 This demonstrates that generating a large number of high-quality data related to the predicted event through GNDiff significantly affects event prediction.
-DPCL leverages Poincaré space properties to distinguish similar periodic events better, enhancing inference accuracy.
+DPCL leverages the properties of Poincaré space to better capture the relationships between periodic entities and distinguish similar periodic events, thereby enhancing inference accuracy.
+Effectiveness Analysis of Graph Diffusion in New Events
+To evaluate the effectiveness of graph diffusion models on new events, we curated datasets from four sources by selecting the first occurrences of distinct events, thereby focusing exclusively on new data.
+We extracted datasets based on the first appearance of all events that occurred, ensuring that only new and unique occurrences were included, while subsequent appearances of the same events were excluded.
+Detailed dataset statistics are provided in Table 7, which supports our methodological approach.
+The performance comparisons, detailed in Table 5 and Table 6, show that the GNDiff model consistently outperforms the DPCL model across all metrics and datasets.
+This robust performance highlights the model’s capability to adeptly handle dynamic, unseen data, proving GNDiff’s exceptional utility in environments that require the continuous integration and processing of new information.
 Method YAGO
@@
 Table 8: Experimental results of DPCL under different spatial mapping strategies
-4.4 Dual Domain Mapping Strategy Analysis
-In this study, we separately examined the effects of different spatial mapping strategies on the learning effect of DPCL.
-Table 4 shows the results of the experiments on the YAGO dataset without the GNDiff.
-Euc/Hyp denotes that periodic and non-periodic event entities are mapped to Euclidean and Poincaré spaces respectively for comparative learning, Hyp/ Euc denotes that the two types of entities are mapped to Poincaré and Euclidean space, Hyp/Hyp denotes that both types of entities are mapped to Poincaré space, and Euc / Euc denotes that the two types of entities are mapped to Euclidean space respectively.
-From the experimental results, it can be seen that the Hyp/Euc method is the most effective, while the Euc/Euc method is the least effective.
-This suggests that mapping periodic and non-periodic event entities to separate spaces can improve the modeling of comparative learning.
-Specifically, the geometric properties of the Poincaré space can effectively distinguish entities whose feature representations are too close to each other in the Euclidean space.
-5 Conclusion
+Dual Domain Mapping Strategy Analysis
+In this study, we examined the effects of different spatial mapping strategies on DPCL’s learning performance.
+Table 8 presents the results on the YAGO dataset without GNDiff.
+Euc/Hyp denotes that periodic and non-periodic event entities are mapped to Euclidean and Poincaré spaces, respectively, for comparative learning, Hyp/ Euc denotes that the two types of entities are mapped to Poincaré and Euclidean space, Hyp/Hyp denotes that both types of entities are mapped to Poincaré space.
+Euc / Euc denotes that the two types of entities are mapped to Euclidean space, respectively.
+The experimental results show that the Hyp/Euc method is the most effective, while the Euc/Euc method is the least effective.
+This suggests that mapping periodic and non-periodic event entities to separate spaces can improve the modelling of comparative learning.
+Specifically, compared to Euclidean space, the geometric properties of the Poincaré space allow for a more accurate capture of the relationships between periodic entities and distinguish similar ones, leading to better link prediction results.
+Dual Domain Embedded Visual Analytics
+In this section, we visualize entity pairs—(Germany, Berlin) and (Germany, Munich)—with the same relation, embedded in both Euclidean and Poincaré spaces, as shown in Fig.
+2.
+In Euclidean space, the relationships between Germany and the two cities are not well differentiated.
+In contrast, Poincaré space provides a clearer distinction, better reflecting the relationships between the entities.
+By leveraging the geometric properties of Poincaré space, we improve the model’s ability to capture entity relationships, better distinguish similar entities, thereby enhancing prediction accuracy.
+Graph Node Diffusion Generation Process
+To address the issue of insufficient interaction traces for new events on the timeline, GNDiff generates data samples by simulating new events using a diffusion mechanism.
+Table 9 illustrates this process for GNDiff.
+Since graph node data differs from text data, we first obtain node embedding vectors and then apply the diffusion mechanism.
+Each node is transformed into a 128-dimensional vector, but for simplicity, we use a 4-dimensional vector to demonstrate the process.
+The process begins with a fully masked state and progressively reveals information about entities and relationships.
+Initially (t = 0), all information is masked.
+For instance, even though entity 1 (United States) and the relation (sanctioned) are known, they are masked, requiring the model to predict entity 2.
+By step three (t = 3), the model predicts entity 2 based on frequency information, generating a vector [0.23, 0.37, 0.52, ?].
+In step four (t = 4), the model adjusts this vector to [0.18, 0.29, 0.71, ?].
+Finally, the model generates the complete vector [0.18, 0.29, 0.71, 0.86], mapping it to the entity ”Russia,” thus completing the generation process.
+This iterative adjustment enhances the quality of samples, improving the model’s accuracy and robustness in predicting new events.
+Hyper-parameter Analysis
+DPCL-Diff involves two key hyperparameters, $\alpha$ and $\lambda$, whose impact on performance is analyzed for the ICEWS14 and YAGO datasets in Fig.
+3.
+$\alpha$ balances $L_{\text{diff}}$ and $L_{\text{dpcl}}$.
+On ICEWS14, $\alpha=0.3$ enhances temporal interaction modeling with minimal sensitivity to changes.
+For YAGO, a lower $\alpha$ is more effective, highlighting the importance of dual-domain contrastive learning for periodic events.
+For $\lambda$, after fixing $\alpha$, $\lambda=4$ optimizes performance on ICEWS14 by balancing temporal variations and preventing overfitting.
+On YAGO, progressively increasing $\lambda$ improves performance, handling the dataset’s complex structure.
+These results demonstrate DPCL-Diff’s robustness and ability to maintain high performance with stable hyperparameter settings, effectively balancing diffusion and contrastive learning across datasets.
+Figure 2: Illustration of 2D entity embedding learned by Euc (left) and Hyp (right) on YAGO.
+step Entity 1 Relation Entity 2
+t = 0 [MASK] [MASK] [MASK]
+t = 1 United States [MASK] [MASK]
+t = 2 United States sanctioned [MASK]
+t = 3 United States sanctioned [0.23, 0.37, 0.52, ?]
+t = 4 United States sanctioned [0.18, 0.29, 0.71, ?]
+t = 4 United States sanctioned [0.18, 0.29, 0.71, 0.86]
+t = 5 United States sanctioned Russia
+Table 9: Example of graphical node diffusion model generation process
+Figure 3: Results of hyper-parameters $\alpha$ and $\lambda$ of DPCL-Diff on ICEWS14 and YAGO.
+Conclusion
 In this paper, we propose a temporal knowledge graph inference method based on a graph node diffusion model with dual-domain periodic contrastive learning.
@@
 Extensive experiments demonstrate the reasoning advantages and effectiveness of DPCL-Diff on public temporal knowledge graph datasets.
-6 Limitations
-This study does not incorporate adaptive embedding strategies to distinguish between periodic and non-periodic events.
-Consequently, our approach lacks the flexibility to dynamically adjust to different types of temporal knowledge graph data.
-This limitation could potentially impact the model’s effectiveness when applied to datasets with varying temporal characteristics.
-Appendices
-7 Extended Experiments
-7.1 Dual Domain embedded visual analytics
-In this section, we investigate the visualization of entities in DPCL with bi-domain embedding.
-As shown in Fig.
-2, we visualize two sets of entity pairs—(Germany, Berlin) and (Germany, Munich)—with the same relation, mapping them to Euclidean and Poincaré space, respectively.
-In Euclidean space, due to the similar Euclidean distance between Berlin and Munich to Germany, it is difficult to distinguish between these two entities effectively.
-However, in Poincaré space, Germany is significantly farther from Munich, bringing Germany closer to Berlin and making it easier to derive the correct inference.
-By leveraging the spatial properties of Poincaré space, we can more effectively distinguish similar periodic event entities and improve the overall prediction accuracy of the model.
-Figure 2: Illustration of 2D entity embedding learned by Euc (left) and Hyp (right) on YAGO.
+Acknowledgments
+This work was supported by the National Natural Science Foundation of China (No. 62402307, No. U23B2021).